OpenStudy (anonymous):

Help please : differentiate cos 2x ? how to solve this ?

OpenStudy (anonymous):

is the answer -2sinx ?

OpenStudy (anonymous):

Close, it's -2sin2x

OpenStudy (anonymous):

Saying there are "two function." Let f(x) be cosx and g(x) be 2x. Apply chain rule: d/dx f(g(x)) = f'(g(x))·g'(x) cos(2x) = -sin(2x) · 2 = -2sin(2x) Is this clear?

OpenStudy (anonymous):

ok i understand how we got -2sinx but how did we get 2x ?

OpenStudy (anonymous):

ahhh ok i got it now !! thanks

OpenStudy (anonymous):

do you mind if i ask you more questions ?

OpenStudy (anonymous):

While applying chain rule, you derivative first function and left second function untouched, then derivative second function.

OpenStudy (anonymous):

You're welcome. Close this question and post another question.